A box open at the top has a rectangular base 200mmx300mm and an altitude of 150mm. If the base and the sides are 10mm thick, find the total surface area of the box.
Answer by Alan3354(69443)

A box open at the top has a rectangular base 200mmx300mm and an altitude of 150mm. If the base and the sides are 10mm thick, find the total surface area of the box.
---------------
A box open at the top has a rectangular base 20 cm x 30cm and an altitude of 15cm. If the base and the sides are 1cm thick, find the total surface area of the box.
--------------
On the outside:
The SA of the bottom is 20*30 = 600 sq cms
2 sides are 15*20 = 300*2 = 600 sq cms
2 sides are 15*30 = 450*2 = 900 sq cms
Outside SA = 600+600+900 = 2100 sq cms
------------
The inside dim's are the OD's minus 2 cms.
--> 18 X 28 X 13
--------------
The bottom is 18*28 = 504 sq cms
2 sides are 13 X 18 = 468 sq cms
2 sides are 28 X 13 = 728 sq cms
---------------
Area of the open top is 1 cm X the difference of the OD base and the ID base
= 600 - 504 = 96 sq cms
-----------------------
Add those.
If you want to convert to sq mms, multiply by 100.